A farmer grows a bushel of wheat and sells it to a miller for 100 dollar. The miller turns the wheat into flour and sells the flour to a baker for 300 dollar. The baker uses the flour to make bread and sells the bread to university for 600 dollar. The university students eat the bread. What is the value added by students?what is GDP in this case?


1
Expert's answer
2022-01-12T09:11:38-0500

The pupils contribute no value to the product since they only consume the bread and do not process it further.

GDP= Final product cost=birr 600 or

Amount of value added by each person= (100- 0)+(300-100)+(600-300)

=100+200+300= birr 600